OpenStudy (anonymous):

Help me with this problem! 4 - ( 2 - 2 ) x8

OpenStudy (anonymous):

start with parenthesis operation first then M D A S....

OpenStudy (kc_kennylau):

Even the calculator in your home can do it.

OpenStudy (kc_kennylau):

google search is also a calculator which can be used.

OpenStudy (kc_kennylau):

Wolfram|Alpha also works.

OpenStudy (anonymous):

Multiplication and Division have the same level ...

OpenStudy (anonymous):

Just do the problem

OpenStudy (anonymous):

after which you can have either Addition or Subtraction (same level) but lower than M & D...

OpenStudy (kc_kennylau):

~~~~~~PEMDAS EXPLAINED~~~~~~~ P for Parentheses. Do the job inside the parentheses first, following the EMDAS order. E for Exponents. You'll learn it later. M and D for Multiplication and Division, whichever comes first is done first. A and S for Addition and Subtraction, whichever comes first is done first.

OpenStudy (anonymous):

all the help that you need is given... now it's all yours... you can solve it...

OpenStudy (anonymous):

Im asking someone to just to the stupid problem for me not write a essay about it.

OpenStudy (anonymous):

you wouldn't learn if we'll do it for you....

OpenStudy (nincompoop):

if we did your homework for you, our names should appear in the diploma, not yours

OpenStudy (kc_kennylau):

exactly

OpenStudy (kc_kennylau):

but here are the steps. So we have: \[4-(2-2)\times8\] P for parentheses. They go first. So now we have \[4-0\times8\] E skipped. M and D for Multiplication and Division. So now we have \[4-8\] A and S for Addition and Subtraction. So we have \[-4\] In a cluster here it goes.\[4-(2-2)\times8\]\[=4-0\times8\]\[=4-8\]\[=-4\]

OpenStudy (nincompoop):

we are here to guide you with the concepts that you can not grasp sometimes this may involve lengthy explanation and solving DIFFERENT problems of the same concept there may be situations when we cannot give other problems to solve, but that is very rare in math

OpenStudy (kc_kennylau):

But if you just copy the steps to your homework without digesting them, you would get a \(0\) in the exam, because you don't know how to the the same type of questions.

OpenStudy (kc_kennylau):

And your dad and mom will be mad at you.

OpenStudy (kc_kennylau):

And they will take you into tutor class.

OpenStudy (kc_kennylau):

And you will have a lot of stuff to do.

OpenStudy (kc_kennylau):

That's it i'm done talking

OpenStudy (kc_kennylau):

@nincompoop your turn :)